# Reconstructing Sepsis Trajectories from Clinical Case Reports using LLMs: the Textual Time Series Corpus for Sepsis

### v1

@@ 2.4 Annotation Quality Review
To add to the quantitative assessment of annotations at the event-time level, we also verified the annotation quality at the case-report level in our annotation quality review. To inspect the quality of the annotations, we had our physician annotator review the annotations of the Sepsis-10 dataset that the LLM models produced. The annotator was asked to rank-order the annotations (as well as the manual annotation file), with rank 1 being the best, and rank 7 being the worst. The annotator also placed the annotations into quality levels of Excellent, Good, Acceptable, and Poor. The guideline for placing the annotations into each quality level was:
@@
Table 2: Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across LLMs for i2m4, sepsis-10, and sepsis-100 datasets.
i2m4 sepsis-10 sepsis-100
Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC
O1-preview 0.76 0.932 22 0.796
DeepSeek R1 0.35 0.702 16 0.751 0.79 0.876 12 0.772 0.91 0.875 2 0.801
L3.3 70B 0.39 0.730 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791
L3.1 70B 0.30 0.688 21 0.738 0.75 0.915 24 0.741 0.76 0.891 5 0.790
Mixtral 8x7B 0.23 0.731 27 0.726 0.68 0.914 72 0.570 0.71 0.860 24 0.737
L3.1 8B 0.34 0.707 61 0.682 0.69 0.801 376 0.615 0.73 0.78 48 0.705
BERT 5-CV 0.704 34 0.520
@@ 3 Results
Our analysis resulted in T2S2, a new open-access textual time series corpus for Sepsis-3, comprising 2,139 case reports. Demographically, the corpus is 58% male, 42% female, with ages ranging from 0 to 111 and a mean age of 49 (IQR: 32-65). The sepsis-10 manual annotations are of 10 files: PMC3075162, PMC4778217, PMC6238558, PMC7576661, PMC9136437, PMC9355071, PMC9552858, PMC10370906, PMC10556716, and PMC10629858, resulting in 801 reference events. An example annotation set (manual, O1-preview, DeepSeek R1, Llama 3.3, Mixtral 8x7B, and Llama 3.1 8B) for the last case is shown in Appendix C.
@@ 3.2 Annotation Results
The results in Table 2 show the strong performance of O1-preview, DeepSeek R1 IQ1, and Llama-3.3-70B-Instruct across all three datasets. In i2m4, all match rates are low (BERT 5-CV is omitted since the method provides the event list (Frattallone-Llado et al., 2024)), highlighting the difficulty of matching short text spans corresponding to “concepts” rather than chaining them at the level of a clinical finding. Among the matches, Llama 3.3 and DeepSeek R1 performed best for time predictions (trading off event matching with temporal performance), and substantially outperforming the baseline BERT model, particularly with respect to time discrepancy. However, the overall results on i2m4 were considerably lower than in the PMOA datasets, including in terms of time ordering and time discrepancy. In contrast to concordances of 0.707 in i2m4, in sepsis-10, O1-preview and Llama 3.3 were excellent time-ordering annotators, both with concordances of 0.932.
Table 3: Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across Llama-3.3-70B-Instruct prompt variants: no role-playing (no role), no conjunction expansion (no expansion), zero-shot, interval, and interval with i2b2 event typing (Interval + Type), for i2m4, sepsis-10, and sepsis-100.
i2m4 sepsis-10 sepsis-100
Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC
L3.3 70B 0.40 0.707 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791
Ablations
No role 0.37 0.670 24 0.738 0.75 0.920 24 0.743 0.74 0.887 4 0.789
No expansion 0.37 0.659 23 0.723 0.73 0.906 24 0.753 0.78 0.886 3 0.793
Zero-shot 0.27 0.632 17 0.741 0.56 0.748 24 0.746 0.69 0.803 6 0.795
Augments
Interval 0.41 0.685 15 0.713 0.77 0.921 48 0.725 0.78 0.887 5 0.791
Interval+Type 0.38 0.653 13 0.695 0.66 0.918 24 0.736 0.76 0.867 6 0.776
To examine the results in greater detail, we show figures of the cumulative distribution function of event matches and concordance boxplots (Figure 2) for the sepsis-10 dataset. When the function reaches its maximum at the maximum cosine distance, it may be less than 1 because the label events have no events left to match, i.e., the reference method recovered an event that the prediction method did not. The event match rates in Table 2 correspond to the match rate at the cosine distance threshold in Figure 2. One could adjust the cosine distance threshold, for example, by increasing it to increase matches, at the cost of false positive matches, which degrade the temporal performance characteristics. The concordance box plots (Figure 2) show strong performance among the large LLMs, whereas the Llama 3.1 8B and Mixtral 8x7B model performances are much lower, which is a consistent finding throughout our analysis.
@@
The sensitivity analysis to our chosen prompt shows the relative value of various elements for the Llama-3.3-70B-Instruct model (Table 3). Our original prompt appears to be the strong performer overall, showcasing the value in particular of few-shot prompting, but also the value of the technical specification of how to treat conjunctions and the modest benefit of role-playing. Surprisingly, when comparing the main L3.3 70B model responses to the augmentations responses, it appears that requesting timestamp intervals (Interval) modestly increases the event match rate. However, the temporal performance characteristic degrades, suggesting a possible request overload on the model or an increase in false positive event matches. The performance of the typing method (Interval + Type) is further degraded. Detailed assessments for the sepsis-10 ablations and the i2m4 dataset, analogous to Figure 2, are shown in Appendix D. To better understand the performance differences across datasets, we conducted an error analysis between sepsis-10 and i2m4, which is presented in the Appendix E. This analysis provides additional insight into the factors contributing to variations in model performance for event match rate and c-index across datasets.
Figure 2: Event match cumulative distribution function (left), concordance box-plots (middle left), time discrepancy from the manual annotation timestamps among matched events, overall (middle right), and disaggregated by clinician annotator timestamp (time from presentation, right) for the sepsis-10 dataset.
@@ 3.3 Annotation Quality Results
Inspecting the clinical quality of the annotations, we report the mean rank, top-1 fraction, top-1-LLM fraction, and quality level in Table 4. There is a clear performance ordering, with O1-preview annotations being a close second to the manual annotations, followed by DeepSeek R1 IQ1 and Llama 3.3. While the Llama 3.3 annotations are found to be acceptable a high percentage of the time (90%), the other Llama and Mistral approaches fail to achieve Acceptable performance in excess of 50%. Comparing the DeepSeek R1 IQ1 and Llama 3.3 annotations, the former has more variable performance (only 8 of 10 with Acceptable annotations), but maintains both higher match rate and AULTC (Table 2 and Figure 2 (left and middle right)). Our clinical reviewer commented that the Llama 3.3 discrepancies for large time horizons, i.e., years, were large, which is corroborated in Figure 2 (right, bottom subpanels).
@@
Events with duration. Example: “10-kg weight loss” from “he was presented to the hospital with complaints of abdominal distension, constipation, vomiting, and a 10-kg weight loss.” This example from case report PMC10629858 references a gradual event that has been ongoing for 2 months. By the annotation guideline for selecting the start time of events with duration, the time -168 (what O1-preview annotates) is correct. However, the event can also be considered as “history of 10-kg weight loss” via the “history of” guideline with time 0 (what L3.3 70B annotates). Note, this example could be resolved by interval representations, but we see the annotation quality (particularly the AULTC) degrade when requesting additional fields (Table 3 and Figure D.1).
@@ 4 Discussion and Conclusions
Table 4: Sepsis-10 annotation quality review
Annotator Mean Rank Top-1 Top-1-LLM Excellent $\geq$ Good $\geq$ Acceptable
Manual 1.2 0.9 1 1 1
O1-preview 1.5 0.6 0.9 0.9 1 1
DeepSeek R1 IQ1 3.5 0.1 0.2 0.4 0.5 0.8
L3.3 70B 3.8 0 0 0 0.2 0.9
L3.1 70B 4.8 0 0 0 0.2 0.5
Mixtral 8x7B 5.4 0 0 0 0.1 0.2
L3.1 8B 6.5 0 0 0 0 0
@@
Regarding generalizability to other medical conditions besides sepsis, our methodology could be extended in a straightforward manner. Our motivation for choosing sepsis as the disease focus was to expand the analysis from the i2m4 dataset to a similar population in the PMOA corpus and maintain a well-defined phenotype. Sepsis is a sensible choice as it is one of the most prevalent and important diagnoses in critical care medicine. However, there are several concerns when generalizing to other conditions. For one, most other conditions are less acute in nature, meaning that the time distributions may be wider. Also, the amount and type of information in such case reports may vary. Despite this, one of the benefits of the AULTC (compared to linear/squared discrepancies) when you mix intensive care and chronic disease cases is that the chronic disease discrepancies at wider times scales will not dominate the performance measure (because of the time discrepancy log scaling). Another generalizability concern is that for some conditions, less temporal information will be present. This appears to be more acute in imaging studies, where the focus of the case is on the image content at a single snapshot in time. For these types of studies, the longitudinal annotations become less meaningful compared to the image’s information content.
Future work could further characterize the findings, with additional or improved augmentations to the current system. Although our augmentation strategies did not improve overall system performance (Table 3), alternative prompting and or fine-tuning frameworks could further improve the annotations in terms of quality and representation. Our hope is that this corpus, or a future version of it, will be used for downstream applications in sepsis, such as causal modeling, forecasting, and multimodal alignment. As a community, we must use the best data available regardless of its format, and for temporal tasks, this temporal, textual approach offers a complementary corpus to those in frequent use. Further developments in temporal extraction could improve upon our pipeline and could be tailored to one or more of the applications above, each of which place different emphasis on what constitutes a high-quality clinical time series. For example, time-ordering is a necessary condition for causality, whereas the precise timestamp may be of high importance in forecasting and alignment. Finally, while our clinical finding and timestamp pairs attempt to capture the true occurrence and timing of each event, another approach focusing on the timing of perceived occurrences could better uncover how actors process and affect their environments, leading to better characterization for decision analysis, multi-agent systems, and resource allocation. Multimodal integrations could help account for perception differences and design collection strategies to assist early detection of sepsis and sepsis progression.

### v2

@@ 2.4 Annotation Quality Review
To add to the quantitative assessment of annotations at the event-time level, we also verified the annotation quality at the case-report level in our annotation quality review. To inspect the quality of the annotations, we had our physician annotator review the annotations of the Sepsis-10 dataset that the LLM models produced. The annotator was asked to rank-order the annotations (as well as the manual annotation file), with rank 1 being the best, and rank 7 being the worst. The annotator also placed the annotations into quality levels of Excellent, Good, Acceptable, and Poor. The guideline for placing the annotations into each quality level:
@@
[…]
Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across LLMs for i2m4, sepsis-10, and sepsis-100 datasets. i2m4 sepsis-10 sepsis-100 Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC O1-preview 0.76 0.932 22 0.796 DeepSeek R1 0.35 0.702 16 0.751 0.79 0.876 12 0.772 0.91 0.875 2 0.801 L3.3 70B 0.39 0.730 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791 L3.1 70B 0.30 0.688 21 0.738 0.75 0.915 24 0.741 0.76 0.891 5 0.790 Mixtral 8x7B 0.23 0.731 27 0.726 0.68 0.914 72 0.570 0.71 0.860 24 0.737 L3.1 8B 0.34 0.707 61 0.682 0.69 0.801 376 0.615 0.73 0.78 48 0.705 BERT 5-CV 0.704 34 0.520
@@ 3 Results
Our analysis resulted in T2S2, a new open-access textual time series corpus for Sepsis-3, comprising 2,139 case reports. Demographically, the corpus is 58% male, 42% female, with ages ranging from 0 to 111 and a mean age of 49 (IQR: 32-65). The sepsis-10 manual annotations are of 10 files: PMC3075162, PMC4778217, PMC6238558, PMC7576661, PMC9136437, PMC9355071, PMC9552858, PMC10370906, PMC10556716, and PMC10629858, resulting in 801 reference events. An example annotation set (manual, O1-preview, DeepSeek R1, Llama 3.3, Mixtral 8x7B, and Llama 3.1 8B) for the last case is shown in Appendix .1.
@@ 3.2 Annotation Results
The results in Table 2.4 show the strong performance of O1-preview, DeepSeek R1 IQ1, and Llama-3.3-70B-Instruct across all three datasets. In i2m4, all match rates are low (BERT 5-CV is omitted since the method provides the event list [9]), highlighting the difficulty of matching short text spans corresponding to “concepts” rather than chaining them at the level of a clinical finding. Among the matches, Llama 3.3 and DeepSeek R1 performed best for time predictions (trading off event matching with temporal performance), and substantially outperforming the baseline BERT model, particularly with respect to time discrepancy. However, the overall results on i2m4 were considerably lower than in the PMOA datasets, including in terms of time ordering and time discrepancy. In contrast to concordances of 0.707 in i2m4, in sepsis-10, O1-preview and Llama 3.3 were excellent time-ordering annotators, both with concordances of 0.932.
[…]
Comparison of event match rate (Event), median concordance ($c$), median absolute error (MAE), and area under the log-time curve (AULTC) across Llama-3.3-70B-Instruct prompt variants: no role-playing (no role), no conjunction expansion (no expansion), zero-shot, interval, and interval with i2b2 event typing (Interval + Type), for our 3 datasets. i2m4 sepsis-10 sepsis-100 Model Event c MAE AULTC Event c MAE AULTC Event c MAE AULTC L3.3 70B 0.40 0.707 27 0.721 0.75 0.932 24 0.759 0.78 0.894 4 0.791 Ablations No role 0.37 0.670 24 0.738 0.75 0.920 24 0.743 0.74 0.887 4 0.789 No expansion 0.37 0.659 23 0.723 0.73 0.906 24 0.753 0.78 0.886 3 0.793 Zero-shot 0.27 0.632 17 0.741 0.56 0.748 24 0.746 0.69 0.803 6 0.795 Augments Interval 0.41 0.685 15 0.713 0.77 0.921 48 0.725 0.78 0.887 5 0.791 Interval+Type 0.38 0.653 13 0.695 0.66 0.918 24 0.736 0.76 0.867 6 0.776
To examine the results in greater detail, we show figures of the cumulative distribution function of event matches and concordance boxplots (Figure 2) for the sepsis-10 dataset. When the function reaches its maximum at the maximum cosine distance, it may be less than 1 because the label events have no events left to match, i.e., the reference method recovered an event the prediction method did not. The event match rates in Table 2.4 correspond to the match rate at the cosine distance threshold in Figure 2. One could adjust the cosine distance threshold, for example, by increasing it to increase matches, at the cost of false positives, which degrade temporal performance characteristics. The concordance boxplots (Figure 2) show strong performance among the large LLMs, whereas the Llama 3.1 8B and Mixtral 8x7B model performances are much lower—a consistent finding throughout our analysis.
@@
The sensitivity analysis to our chosen prompt shows the relative value of various elements for the Llama-3.3-70B-Instruct model (Table 3.2). Our original prompt appears to be the strong performer overall, showcasing the value in particular of few-shot prompting, but also the value of the technical specification of how to treat conjunctions and the modest benefit of role-playing. Surprisingly, when comparing the main L3.3 70B model responses to the augmentations responses, it appears that requesting timestamp intervals (Interval) modestly increases the event match rate. However, the temporal performance characteristic degrades, suggesting a possible request overload on the model or an increase in false positive event matches. The performance of the typing method (Interval + Type) is further degraded. Detailed assessments for the sepsis-10 ablations and the i2m4 dataset, analogous to Figure 2, are shown in Appendix Reconstructing Sepsis Trajectories from Clinical Case Reports using LLMs: the Textual Time Series Corpus for Sepsis. To better understand the performance differences across datasets, we conducted an error analysis between sepsis-10 and i2m4, which is presented in the Appendix .5. This analysis provides additional insight into the factors contributing to variations in model performance for event match rate and c-index across datasets.
Figure 2: Event match CDF (left), concordance box-plots (middle left), time discrepancy from the manual annotation timestamps among matched events, overall (middle right), and disaggregated by clinician annotator timestamp (time from presentation, right) for the sepsis-10 dataset.
@@ 3.3 Annotation Quality Results
Inspecting the clinical quality of the annotations, we report the mean rank, top-1 fraction, top-1-LLM fraction, and quality level in Table 4. There is a clear performance ordering, with O1-preview annotations being a close second to the manual annotations, followed by DeepSeek R1 IQ1 and Llama 3.3. While the Llama 3.3 annotations are found to be acceptable a high percentage of the time (90%), the other Llama and Mistral approaches fail to achieve Acceptable performance in excess of 50%. Comparing the DeepSeek R1 IQ1 and Llama 3.3 annotations, the former has more variable performance (only 8 of 10 with Acceptable annotations), but maintains both higher match rate and AULTC (Table 2.4 and Figure 2 (left and middle right)). Our clinical reviewer commented that the Llama 3.3 discrepancies for large time horizons, i.e., years, were large, which is corroborated in Figure 2 (right, bottom subpanels).
@@
Events with duration. Example: “10-kg weight loss” from “he was presented to the hospital with complaints of abdominal distension, constipation, vomiting, and a 10-kg weight loss.” This example from case report PMC10629858 references a gradual event that has been ongoing for 2 months. By the annotation guideline for selecting the start time of events with duration, the time -168 (what O1-preview annotates) is correct. However, the event can also be considered as “history of 10-kg weight loss” via the “history of” guideline with time 0 (what L3.3 70B annotates). Note, this example could be resolved by interval representations, but we see the annotation quality (particularly the AULTC) degrade when requesting additional fields (Table 3.2 and Figure .3).
@@ 4 Discussion and Conclusions
[…]
Sepsis-10 annotation quality review Annotator Mean Rank Top-1 Top-1-LLM Excellent $\geq$ Good $\geq$ Acceptable Manual 1.2 0.9 1 1 1 O1-preview 1.5 0.6 0.9 0.9 1 1 DeepSeek R1 IQ1 3.5 0.1 0.2 0.4 0.5 0.8 L3.3 70B 3.8 0 0 0 0.2 0.9 L3.1 70B 4.8 0 0 0 0.2 0.5 Mixtral 8x7B 5.4 0 0 0 0.1 0.2 L3.1 8B 6.5 0 0 0 0 0
@@
Regarding generalizability to other medical conditions besides sepsis, our methodology could be extended in a straightforward manner. Our motivation for choosing sepsis as the disease focus was to expand the analysis from the i2m4 dataset to a similar population in the PMOA corpus and maintain a well-defined phenotype. Sepsis is a sensible choice as it is one of the most prevalent and important diagnoses in critical care medicine. However, there are several concerns when generalizing to other conditions. For one, most other conditions are less acute in nature, meaning that the time distributions may be wider. Also, the amount and type of information in such case reports may vary. Despite this, one of the benefits of the AULTC (compared to linear/squared discrepancies) when mixing intensive care and chronic disease cases is that the chronic disease discrepancies at wider times scales do not dominate the performance measure (because of the time discrepancy log scaling).
Another generalizability concern is that for some conditions, less temporal information will be present. This appears to be more acute in imaging studies, where the focus of the case is on the image content at a single snapshot in time. For these types of studies, the longitudinal annotations become less meaningful compared to the image’s information content.
Future work could further characterize the findings, with additional or improved augmentations to the current system. Although our augmentation strategies did not improve overall system performance (Table 3.2), alternative prompting and or fine-tuning frameworks could further improve the annotations in terms of quality and representation. Our hope is that this corpus, or a future version of it, will be used for downstream applications in sepsis, such as causal modeling, forecasting, and multimodal alignment. As a community, we must use the best data available regardless of its format, and for temporal tasks, this temporal, textual approach offers a complementary corpus to those in frequent use. Further developments in temporal extraction could improve upon our pipeline and could be tailored to one or more of the applications above, each of which place different emphasis on what constitutes a high-quality clinical time series. For example, time-ordering is a necessary condition for causality, whereas the precise timestamp may be of high importance in forecasting and alignment. Finally, while our clinical finding and timestamp pairs attempt to capture the true occurrence and timing of each event, another approach focusing on the timing of perceived occurrences could better uncover how actors process and affect their environments, leading to better characterization for decision analysis, multi-agent systems, and resource allocation. Multimodal integrations could help account for perception differences and design collection strategies to assist early detection of sepsis and sepsis progression.